# The SRG/eROSITA all-sky survey

### v1

@@ Abstract
Context. The Fornax cluster is one of the most nearby X-ray bright galaxy clusters, allowing us to study the system at high spatial resolution. However, previous observations of the intracluster medium were limited to less than .
@@
Methods. We use data from five SRG/eROSITA all-sky surveys and perform a detailed one- and two-dimensional X-ray surface brightness analysis, allowing us to trace hot gas emission from kpc to Mpc scales with a single instrument. We compare the results to those from a recent numerical simulation of the local Universe (SLOW) and correlate the X-ray emission distribution with that of other tracers, including cluster member galaxies, ultra compact dwarf galaxies, intracluster globular clusters, and HI-tail galaxies.
Results. We detect X-ray emission out to well beyond the virial radius, deg. We do not find obvious evidence for the bow shock several hundred kpc south of the cluster center predicted by previous numerical simulations of the Fornax cluster. Instead, we discover finger-like structures beyond to the west and south-east that stretch out far beyond the virial radius. They might be due to gas being pushed outward by the previous merger with NGC 1404 or due to warm-hot gas infall along large-scale filaments. Intriguingly, we find the distributions of the other tracers – galaxies and globular clusters – to be correlated with the X-ray excess regions, favoring the latter infall scenario. Interestingly, we also discover an apparent Bridge of low surface brightness emission beyond the virial radius connecting to the Fornax A galaxy group, which is also traced by the member galaxy and globular cluster distribution. This X-ray Bridge furthermore approximately coincides with a region of enhanced Faraday depth detected previously. The gas distribution in the SLOW simulation shows similar features as those we have discovered with SRG/eROSITA.
Conclusions. SRG/eROSITA has allowed us to tremendously expand the view of the intracluster medium of the Fornax cluster. We witness the growth of a cluster along large-scale filaments.
@@
With a virial diameter of almost 4.5 deg, the Fornax cluster is one of the most extended clusters of galaxies in the sky. In fact, only the Virgo cluster (e.g., McCall et al., 2024) spans a larger area. The apparent large extent is due to their proximity. Their intrinsic physical sizes, i.e., masses, are actually not extraordinary; in fact, the Fornax cluster should rather be called a group of galaxies based on its mass of (e.g., Jones et al., 1997; Drinkwater et al., 2001; Reiprich & Böhringer, 2002; Piffaretti et al., 2011; Smith Castelli et al., 2024). Given its proximity, the Fornax cluster can be studied at high spatial resolution and down to low surface brightness at all wavelengths. And, indeed, it is one of the best studied galaxy clusters in existence. In particular, its two central bright elliptical galaxies, NGC 1399 and NGC 1404, as well as the central galaxy of the neighboring Fornax A galaxy group, NGC 1316, have been the subject of many previous studies.
For example, Su et al. (2017b) used a mosaic of 15 XMM-Newton observations (see also Murakami et al. 2011) and Chandra observations to describe four contact discontinuities identified as sloshing cold fronts (for a review about cold fronts see, e.g., Markevitch & Vikhlinin, 2007). They suggest these cold fronts are caused by the infall of NGC 1404.
@@
Deep Chandra data have been used to study NGC 1404 at high spatial resolution, resulting in the findings of a sharp contact discontinuity to the north-west, i.e., in the direction of NGC 1399 at the center of the Fornax cluster, and a tail to the south-east (e.g., Machacek et al., 2005; Su et al., 2017a), consistent with NGC 1404 falling into the core of the Fornax cluster as already indicated by ROSAT PSPC data (Jones et al., 1997).
@@
Sheardown et al. (2018) then used detailed hydrodynamic simulations to study if a presumed infall of NGC 1404 could give rise to the central ICM morphology of the Fornax cluster, e.g., the cold fronts described above as well as the contact discontinuity and tail of NGC 1404. And, indeed, they found the most likely scenario to be that NGC 1404 is on its second or third encounter with the Fornax center. Furthermore, they predicted a detached bow shock 450 kpc to 750 kpc south of NGC 1404. With the eROSITA data used here, we put this prediction to the test.
@@
NGC 1316 is the central galaxy of a group of galaxies that, based on galaxy dynamics, appears to be bound to the Fornax cluster and infalling to it for the first time (Drinkwater et al., 2001). This group is dominated by late-type galaxies and its low fraction of intra-group light points to an ongoing phase of preprocessing (Raj et al., 2020). NGC 1316 is also host to the well-known Fornax A extended radio source and its radio morphology is consistent with a northward, i.e., infalling motion (e.g., Ekers et al., 1983). Recently, Anderson et al. (2021) studied the Faraday depth in a large field surrounding the Fornax cluster using a large number of background sources (see also Loi et al. 2025 for a smaller but more densely sampled Faraday Rotation Measure grid). Anderson et al. found a region of increased Faraday depth north-east of NGC 1316, i.e., in the direction of NGC 1399. In the light of these findings, it is of interest to search for warm-hot gas emission beyond the virial radius of the Fornax cluster towards the direction of Fornax A, which is what we do with the eROSITA data analyzed here.
@@
Given their higher velocity dispersion compared to big galaxies, Drinkwater et al. (2001) concluded that the dwarf galaxy population includes many infalling objects. Indeed, the spatial distribution of dwarf galaxies in the main Fornax cluster is highly structured along clumpy overdensities (e.g., Ordenes-Briceño et al., 2018). The distribution of dwarf galaxies in the infalling Fornax A group is lopsided towards the main cluster, further supporting recent interaction between both components (Su et al., 2021). Given these findings, it is, therefore, of interest to compare the spatial distribution of dwarf galaxies with regions of excess X-ray emission and the warm-hot gas distribution in the cluster outskirts, which may indicate preferred filamentary inflow directions.
@@
Furthermore, we also compare these to the globular cluster and normal galaxy distributions. The inhomogeneous and partly patchy spatial distribution of the intracluster light (Iodice et al., 2016, 2017) and globular clusters (D’Abrusco et al., 2016; Cantiello et al., 2020) within the virial radius trace the recent assembly history of the Fornax cluster. Kinematically cold substructures of intracluster globular clusters (Schuberth et al., 2008; Chaturvedi et al., 2022; Napolitano et al., 2022) support the assembly of the cluster along filaments.
@@
The X-ray telescope eROSITA (Predehl et al., 2021) onboard the Spectrum-Roentgen-Gamma (SRG) satellite (Sunyaev et al., 2021) has performed four all-sky surveys up to now plus one partial sky survey. The data of the first all-sky survey and the corresponding X-ray source catalog is published (Merloni et al., 2024). Given the unlimited field-of-view, these data are perfectly suited to study the very extended Fornax cluster and its surroundings. In this paper, we use data from five surveys.
@@
According to the MCXC database (Piffaretti et al., 2011) the Fornax cluster is located at R.A. 54.6163 and Dec. 35.4483 (J2000) at a redshift of 0.005 and has deg kpc. is the radius within which the mean total density equals 500 times the critical density. Using the conversion in Reiprich et al. (2013), this implies deg kpc and deg kpc; i.e., the virial region of the Fornax cluster spans almost 4.5 deg across. The Fornax cluster lies projected in a region of the sky that features a rather simple X-ray background structure (Fig. 1)111https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/HalfSkyMaps_dr1/ allowing us to perform a robust background treatment despite the large extent.
@@
For the flat cosmology with and km/s/Mpc, assumed throughout this paper, and for a redshift of 0.005, 1 arcsec corresponds to 0.103 kpc.
@@
Since we are interested in the faintest cluster outskirt emission regions, we take advantage of eROSITA data from all five currently existing sky surveys, based on processing version c020. For the data reduction, we use eSASS (Brunner et al., 2022) version users_211214.0.4.
@@
We follow closely the steps described in detail in Reiprich et al. (2021); Veronica et al. (2024, 2025); McCall et al. (2024); Dietl et al. (2024). Therefore, we do not repeat the description of these steps here but only briefly outline them.
@@
We employ a sufficient number of tiles around the Fornax cluster center to cover area at least out to in all directions. We use data from all seven telescope modules (TMs) and apply the parameters pattern=15 and flag=0xe00fff30 when running evtool. We use the latter strict flag filtering since we are mostly interested in the faintest regions, requiring the cleanest data.222See https://erosita.mpe.mpg.de/edr/DataAnalysis/prod_descript/EventFiles_edr.html for details on flag settings. We run flaregti using photons with energy above 5 keV and setting a threshold for the lightcurve filtering. Then, images and exposure maps are created in the energy band 0.2–2.3 keV (0.8–2.3 keV) for the TMs with (without) on-chip filter (Predehl et al., 2021). To generate false color (RGB) images, also images and exposure maps in additional bands are created. The final images are fully corrected: (i) the particle-induced background (PIB; Freyberg et al., 2020; Yeung et al., 2023) is subtracted; (ii) a relative correction for the varying foreground absorption in the field is performed using data from HI4PI Collaboration et al. (2016); (iii) the different TM sensitivities and energy bands are taken into account for the total exposure correction, including vignetting.
@@
Since we are primarily interested in very low surface brightness regions, we perform wavelet-filtering to reduce the noise in the final images as described in Reiprich et al. (2021), using an implementation by F. Pacaud of the filtering method in the MR/1 software (Starck & Pierre, 1998). For some images, we excise sources detected with SExtractor (Bertin & Arnouts, 1996) on the wavelet-filtered images, exploiting the elliptical shape information. Figure 2 shows two image versions resulting from this process.
@@
In addition to qualitative imaging analyses we perform quantitative X-ray surface brightness analyses. We generate exposure- and relative absorption-corrected, PIB-subtracted, and compact source-excised surface brightness profiles in various sectors out to . For the profile analysis, we decrease the region of the SExtractor-detected source at the center of the Fornax cluster/NGC 1399 to a circle with radius 30′′ to remove mostly any potential emission from an active nucleus. The level and uncertainty of the X-ray background (CXB) is determined by placing eleven boxes of size 1 deg2 outside and calculating the average surface brightness and its standard deviation. We use the standard deviation instead of the (smaller) statistical uncertainty based on the number of detected events in order to capture the uncertainty induced by large-scale CXB variations, e.g., due to varying diffuse background emission in our Milky Way. The PIB level in the used 0.2(0.8)–2.3 keV energy band is always well below the CXB level. Two example surface brightness profiles for the sectors 0–180 deg and 180–360 deg, covering scales from 3 kpc to 1 Mpc, are shown in Fig. 3
@@ 3.1 X-ray surface brightness distribution
Focussing first on the core region of the Fornax cluster, we show in Fig. 4 the eROSITA surface brightness overlaid onto an optical color image obtained from the Digitized Sky Survey 2.333https://www.eso.org/public/images/eso0949m/ The X-ray emission peak is centered on NGC 1399. Several other of the bright Fornax member galaxies show X-ray emission, the brightest one being NGC 1404. Furthermore, we observe characteristic features that could be due to sloshing as described in Sect. 1.
@@
To highlight these features more clearly, we show in Fig. 5 a false color wavelet-filtered eROSITA RGB image based on individual fully corrected images in the energy bands 0.2(0.8)–1.0 keV, 1.0–1.4 keV, and 1.4–4.0 keV. The hot gas emission appearing in green sticks out in this representation compared to the large number of point sources (mostly active galactic nuclei, AGN, in the background but also some foreground stars), which often appear redder or bluer. Furthermore, in the inset we show a zoom into the core region using all the useful XMM-Newton data available (1.6 Ms total flare-free exposure time of archival observations). The fully corrected and adaptively smoothed XMM-Newton RGB image is constructed in the energy bands: red: 0.3–0.7 keV, green: 0.7–1.5 keV, blue: 1.5–4.0 keV. The same features and sources are visible in both images with XMM-Newton showing more detail and sharpness in the innermost regions and eROSITA extending the field-of-view to beyond .
@@
To further enhance the visualization of the position of the large-scale sloshing cold front (“Edge”) we employ the Gaussian gradient magnitude (GGM) filtering technique (Sanders et al., 2016, as implemented in SciPy, Virtanen et al. 2020) to the eROSITA image and show the result in Fig. 6.
@@
The real power of eROSITA compared to all other X-ray telescopes in existence comes in through its sensitivity to large-scale low surface brightness features. Already the bottom image in Fig. 2 revealed the vast extent of warm-hot gas emission, dwarfing the emission visible in the huge XMM-Newton mosaic image shown in Fig. 5. Note that we do not expect the surface brightness fluctuations in the bottom image in Fig. 2 to be due to photon noise because of the wavelet-filtering process. Visual support for this expectation comes from comparing the eRASS:2 and eRASS3+eRASS4+eRASS5 images in Fig. 19 in the appendix, which both show essentially the same features while coming from completely independent datasets. The quantitative analysis of the surface brightness profiles in different sectors follows below.
@@
To make additional use of eROSITA’s energy resolution, we show in Fig. 7 an RGB image optimized to show low surface brightness fluctuations. Hot group and cluster gas emission is revealed in this image not only through its extent but additionally through its whitish color. Both images, Figs. 2 (bottom) and 7 demonstrate that the Fornax cluster emission extends not only beyond , already unreachable by the XMM-Newton mosaic, but beyond and actually beyond even the virial radius, , in some directions. In these images, we discover finger-like emission extensions beyond in particular towards the western and south-eastern directions (labeled “Finger 1, 2, and 3” in Fig. 2, bottom). These extensions appear to stretch out to well-beyond , in particular “Excess 1 and 2.”
@@
On the other hand, we do not see any obvious evidence in the X-ray surface brightness maps for the shock front predicted by Sheardown et al. (2018) south of NGC 1404. Below, we check quantitatively for features in the profiles in various sectors.
@@
In principle, varying diffuse emission in our Milky Way or systematic uncertainties in the relative absorption correction can result in apparent surface brightness variations. Both effects would mostly be important at the softest energies keV. Therefore, we construct a hard band image (shown in Fig. 20 in the appendix) and check if the emission features discussed above are robust. And, indeed, they are, lending confidence that they are not due to local effects.
@@
The plots in Fig. 8 illustrate examples of surface brightness features discussed above. The uppermost plot simply shows how well eROSITA can recover already known features in the inner regions; i.e., a strong drop of emission just beyond – the “Edge” seen in Fig. 6. The next plot makes clear that “Excess 1 and 2” in the west-north-west beyond discussed above have, indeed, a significantly higher surface brightness than the CXB level as there are many data points consistently above this level. Furthermore, it shows that our CXB level estimate is conservative because the surface brightness of the dark blue regions is even below our CXB level estimate. The next two plots show how the surface brightness drops abruptly and significantly at the edges of Finger 1 and 2; Finger 1 at and Finger 2 between and .
@@
We can now also search for the detached bow shock predicted by Sheardown et al. (2018) to lie 450–750 kpc south of NGC 1399; i.e., roughly in the range between (393 kpc, 1.06 deg) and (816 kpc, 2.22 deg). Since it is not clear where exactly such a shock would be located and given the complexity of the two-dimensional emission in the outskirts, we start with a large sector and then look a smaller ones in different directions. Figure 3 (bottom) shows the southern 180 deg sector. No obvious surface brightness jumps that might be associated with a shock are obvious in the relevant radial range. Figures 9 and 10 show several more sectors. No obvious jump is present in the range 225–315 deg. In the 240–300 deg sector, there appear to be two peaks of emission surrounding . However, consulting the surface brightness profiles of other sectors, in particular 225–270 deg and 240–270 deg, as well as the Finger 3 towards the south-east in Fig. 2, it becomes clear that these are two extended (note that each individual bin covers 60′′; i.e., twice the eROSITA PSF) clumps of gas, not related to a shock front. On the other hand, the sector 180–270 deg does show a very weak indication of a surface brightness jump between and ; however, we consider it too weak, and the overall two-dimensional emission pattern too complex, to indicate a shock.

### v2

@@ Abstract
Context. The Fornax cluster is one of the closest X-ray-bright galaxy clusters; as such, we can study the system at high spatial resolution. However, previous observations of the intracluster medium were limited to less than .
@@
Methods. We used data from five SRG/eROSITA all-sky surveys and performed a detailed one- and two-dimensional X-ray surface brightness analysis, tracing hot gas emission from kiloparsec to megaparsec scales with a single instrument. We compared the results to those from a recent numerical simulation of the local Universe (SLOW) and correlated the X-ray emission distribution with that of other tracers, including cluster member galaxies, ultra-compact dwarf galaxies, intracluster globular clusters, and HI-tail galaxies.
Results. We detect X-ray emission out to well beyond the virial radius, deg. In the inner regions within , we see previously known features, such as a large-scale spiral-shaped edge; however, we do not find obvious evidence of the bow shock several hundred kiloparsecs south of the cluster center predicted by previous numerical simulations of the Fornax cluster. Instead, we discover emission fingers beyond to the west and southeast and excesses that stretch out far beyond the virial radius. They might be due to gas being pushed outward by the previous merger with NGC 1404 or due to warm-hot gas infall along large-scale filaments. Intriguingly, we find the distributions of the other tracers – galaxies and globular clusters – to be correlated with the X-ray-excess regions, favoring the infall scenario. Interestingly, we also discover an apparent bridge of low-surface-brightness emission beyond the virial radius connecting to the Fornax A galaxy group, which is also traced by the member galaxy and globular cluster distribution. This X-ray bridge furthermore approximately coincides with a region of enhanced Faraday depth detected previously. The gas distribution in the SLOW simulation shows similar features as those we have discovered with SRG/eROSITA.
Conclusions. SRG/eROSITA has enabled us to tremendously expand the view of the intracluster medium of the Fornax cluster. We witness the growth of a cluster along large-scale filaments.
@@
Massive galaxy clusters are the youngest objects in the Universe. They continue to form and grow even today by merging with other clusters and via continuous accretion along large-scale filaments connecting clusters. Cluster outskirts around the virial radius trace this growth and the filament connection (e.g., Brüggen et al., 2011; Reiprich et al., 2013; Walker et al., 2019). The intergalactic medium in clusters – the intracluster medium (ICM) – and in filaments – the warm-hot intergalactic medium (WHIM) – reaches X-ray emitting temperatures above K (e.g., Martizzi et al., 2019). The ICM in the inner parts of clusters has been studied for several decades, while their outskirts became accessible less than 20 years ago (see the review articles above and, e.g., the recent studies of the Virgo and Centaurus clusters, McCall et al. 2024; Veronica et al. 2025). The X-ray emission from the WHIM has been discovered only recently. For instance, WHIM filament discoveries around individual nearby cluster systems have been reported by Reiprich et al. (2021) and Dietl et al. (2024) and through statistical stacking analyses by, for example, Zhang et al. (2024). The nearby Fornax cluster is therefore an obvious candidate for a cluster outskirts study and filament search.
@@
With a virial diameter of almost 4.5 deg, the Fornax cluster is one of the most extended clusters of galaxies in the sky. In fact, only the Virgo cluster (e.g., McCall et al., 2024) spans a larger area. The apparent large extent is due to their proximity. Their intrinsic physical sizes, i.e., masses, are actually not extraordinary; in fact, the Fornax cluster should rather be called a group of galaxies based on its mass of (e.g., Jones et al., 1997; Drinkwater et al., 2001; Reiprich & Böhringer, 2002; Piffaretti et al., 2011; Smith Castelli et al., 2024). Given its proximity, the Fornax cluster can be studied at high spatial resolution and down to low surface brightnesses at all wavelengths. And, indeed, it is one of the best studied galaxy clusters in existence. In particular, its two central bright elliptical galaxies, NGC 1399 and NGC 1404, as well as the central galaxy of the neighboring Fornax A galaxy group, NGC 1316, have been the subject of many previous studies.
@@
For example, Su et al. (2017b) used a mosaic of 15 XMM-Newton observations (see also Murakami et al. 2011) and Chandra observations to describe four contact discontinuities identified as sloshing cold fronts (for a review of cold fronts, see, e.g., Markevitch & Vikhlinin, 2007). They suggest these cold fronts are caused by the infall of NGC 1404. Deep Chandra data have been used to study NGC 1404 at high spatial resolution, resulting in the findings of a sharp contact discontinuity to the northwest, i.e., in the direction of NGC 1399 at the center of the Fornax cluster, and a tail to the southeast (e.g., Machacek et al., 2005; Su et al., 2017a), consistent with NGC 1404 falling into the core of the Fornax cluster as already indicated by ROSAT PSPC data (Jones et al., 1997).
@@
Sheardown et al. (2018) then used detailed hydrodynamic simulations to determine whether a presumed infall of NGC 1404 could give rise to the central ICM morphology of the Fornax cluster, for example the cold fronts described above as well as the contact discontinuity and tail of NGC 1404. And, indeed, they found the most likely scenario to be that NGC 1404 is on its second or third encounter with the Fornax center. Furthermore, they predicted a detached bow shock 450 kpc to 750 kpc south of NGC 1404. With the eROSITA data used here, we put this prediction to the test.
@@
NGC 1316 is the central galaxy of a group of galaxies that, based on galaxy dynamics, appears to be bound to the Fornax cluster and infalling onto it for the first time (Drinkwater et al., 2001). This group is dominated by late-type galaxies, and its low fraction of intragroup light points to an ongoing phase of preprocessing (Raj et al., 2020). NGC 1316 also hosts the well-known Fornax A extended radio source, and its radio morphology is consistent with a northward, i.e., infalling, motion (e.g., Ekers et al., 1983). Recently, Anderson et al. (2021) studied the Faraday depth in a large field surrounding the Fornax cluster using a large number of background sources (see also Loi et al. 2025 for a smaller but more densely sampled Faraday rotation measure grid). Anderson et al. found a region of increased Faraday depth northeast of NGC 1316, i.e., in the direction of NGC 1399. In light of these findings, it is of interest to search for warm-hot gas emission beyond the virial radius of the Fornax cluster in the direction of Fornax A, which is what we do with the eROSITA data analyzed here.
@@
Given the higher velocity dispersion compared to big galaxies, Drinkwater et al. (2001) concluded that the dwarf galaxy population includes many infalling objects. The spatial distribution of dwarf galaxies in the main Fornax cluster is highly structured along clumpy overdensities (e.g., Ordenes-Briceño et al., 2018). The distribution of dwarf galaxies in the infalling Fornax A group is lopsided toward the main cluster, further evidence of a recent interaction between the two components (Su et al., 2021). Given these findings, it is of interest to compare the spatial distribution of dwarf galaxies with regions of excess X-ray emission and the warm-hot gas distribution in the cluster outskirts, which may indicate preferred filamentary inflow directions.
@@
We also compared these distributions to the globular cluster and normal galaxy distributions. The inhomogeneous and partly patchy spatial distributions of the intracluster light (Iodice et al., 2016, 2017) and globular clusters (D’Abrusco et al., 2016; Cantiello et al., 2020) within the virial radius trace the recent assembly history of the Fornax cluster. Kinematically cold substructures of intracluster globular clusters (Schuberth et al., 2008; Chaturvedi et al., 2022; Napolitano et al., 2022) are evidence in favor of the assembly of the cluster along filaments.
@@
The X-ray telescope eROSITA (Predehl et al., 2021) on board the Spectrum-Roentgen-Gamma (SRG) satellite (Sunyaev et al., 2021) has performed four all-sky surveys to date, plus one partial sky survey (Coutinho et al., 2022). The data from the first all-sky survey and the corresponding X-ray source catalog are published (Merloni et al., 2024). Given the unlimited field of view, these data are perfectly suited to studying the very extended Fornax cluster and its surroundings. In this paper, we use data from five surveys. The four additional surveys are each of very similar quality as the first survey. Some minor changes include a decreased particle-induced background (PIB; Yeung et al., 2023), an increased frequency of flares, and some additional bad pixels for telescope module (TM) 4 due to micrometeoroid impact (Freyberg et al., 2022).
@@
According to the MCXC database (Piffaretti et al., 2011), the Fornax cluster is located at RA 54.6163 and Dec. 35.4483 (J2000) at a redshift of 0.005 and has deg kpc. is the radius within which the mean total density equals 500 times the critical density. Using the conversion in Reiprich et al. (2013), this implies deg kpc and deg kpc; i.e., the virial region of the Fornax cluster spans almost 4.5 deg across. The Fornax cluster lies projected in a region of the sky that features a rather simple X-ray background structure (Fig. 1)111https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/HalfSkyMaps_dr1/, enabling us to perform a robust background treatment despite the large extent.
@@
Throughout this paper we assume a flat cosmology with and km/s/Mpc. For a redshift of 0.005, 1 arcsec corresponds to 0.103 kpc.
@@
Since we were interested in the faintest emission regions of the cluster outskirts, we took advantage of eROSITA data from all five currently existing sky surveys, based on processing version c020. For the data reduction, we used the eROSITA Science Analysis Software System (eSASS, Brunner et al., 2022) version users_211214.0.4.
@@
We closely followed the steps described in detail in Reiprich et al. (2021), Veronica et al. (2024, 2025), McCall et al. (2024), and Dietl et al. (2024). Therefore, here we only briefly outline these steps.
@@
We employed a sufficient number of tiles around the Fornax cluster center to cover an area at least out to in all directions. We used data from all seven TMs and applied the parameters pattern=15 and flag=0xe00fff30 when running evtool. We used the latter strict flag filtering since we were mostly interested in the faintest regions and therefore required the cleanest data.222See https://erosita.mpe.mpg.de/edr/DataAnalysis/prod_descript/EventFiles_edr.html for details on flag settings. We ran flaregti using photons with energy above 5 keV and setting a threshold for the light curve filtering. Images and vignetted exposure maps were then created in the energy band 0.2–2.3 keV (0.8–2.3 keV) for the TMs with (without) on-chip filter (Predehl et al., 2021). The PIB was assumed constant across the detector (Freyberg et al., 2020), and we modeled the PIB map of each TM as a rescaled version of its un-vignetted exposure map. The PIB normalizations were estimated from the observed hard band 6.0–9.0 keV counts, rescaled to the appropriate energy range using the spectral distribution of PIB measured in the filter-wheel-closed data (Yeung et al., 2023). To generate false color (RGB) images, images and exposure maps in additional bands were also created. The final images were fully corrected: (i) the PIB was subtracted; (ii) a relative correction for the varying foreground absorption in the field was performed using data from HI4PI Collaboration et al. (2016); and (iii) the different TM sensitivities and energy bands were taken into account for the total exposure correction, including vignetting.
@@
Since we were primarily interested in very low surface brightness regions, we performed wavelet-filtering to reduce the noise in the final images as described in Reiprich et al. (2021), using an implementation by F. Pacaud of the filtering method in the MR/1 software (Starck & Pierre, 1998). For some images, we excised sources detected with SExtractor (Bertin & Arnouts, 1996) on the wavelet-filtered images, exploiting the elliptical shape information. Figure 2 shows two image versions resulting from this process (the labeled features are described in detail in Sect. 3).
@@
In addition to qualitative imaging analyses, we performed quantitative X-ray surface brightness analyses. We generated exposure- and relative absorption-corrected, PIB-subtracted, and compact source-excised surface brightness profiles in various sectors out to . For the profile analysis, we decreased the region of the SExtractor-detected source at the center of the Fornax cluster (NGC 1399) to a circle with radius 30′′ to remove mostly any potential emission from an active nucleus. The level and uncertainty of the cosmic X-ray background (CXB) was determined by placing 11 boxes of size 1 deg2 outside and calculating the average surface brightness and its standard deviation. We used the standard deviation instead of the (smaller) statistical uncertainty based on the number of detected events in order to capture the uncertainty induced by large-scale CXB variations, for example due to varying diffuse background emission in our Milky Way. The PIB level in the used 0.2(0.8)–2.3 keV energy band is always well below the CXB level. Two example surface brightness profiles for the sectors 0–180 deg and 180–360 deg, covering scales from 3 kpc to 1 Mpc, are shown in Fig. 3. Given the large sectors, we used bin widths to exploit the full spatial resolution offered by eROSITA.
@@ 3.1 X-ray surface brightness distribution
Focusing first on the core region of the Fornax cluster, we show in Fig. 4 the eROSITA surface brightness overlaid onto an optical color image obtained from the Digitized Sky Survey 2.333https://www.eso.org/public/images/eso0949m/ The X-ray emission peak is centered on NGC 1399. Several other of the bright Fornax member galaxies show X-ray emission, the brightest one being NGC 1404. Furthermore, we observe characteristic features that could be due to sloshing as described in Sect. 1.
@@
To highlight these features more clearly, we show in Fig. 5 (right) a false color wavelet-filtered eROSITA RGB image based on individual fully corrected images in the energy bands 0.2(0.8)–1.0 keV, 1.0–1.4 keV, and 1.4–4.0 keV. The hot gas emission appearing in green sticks out in this representation compared to the large number of point sources (mostly active galactic nuclei in the background but also some foreground stars), which often appear redder or bluer. Furthermore, on the left we show a zoom into the core region using all the useful XMM-Newton data available (1.6 Ms total flare-free exposure time of archival observations). The fully corrected and adaptively smoothed XMM-Newton RGB image is constructed in the energy bands: red: 0.3–0.7 keV, green: 0.7–1.5 keV, blue: 1.5–4.0 keV. The same features and sources are visible in the two images, with XMM-Newton showing more detail and sharpness in the innermost regions and eROSITA extending the field of view to beyond .
@@
To further enhance the visualization of the position of the large-scale sloshing cold front (the “edge”), we applied the Gaussian gradient magnitude (GGM) filtering technique (Sanders et al., 2016, as implemented in SciPy, Virtanen et al. 2020), to the eROSITA image and show the result in Fig. 6. The similarity to the appearance of the edge in the 1.6 Ms XMM-Newton image in Fig. 5 is striking.
Based on spectral analysis, Su et al. (2017b) identified the edge and the overall spiral pattern toward the inner regions as multiple sloshing cold fronts because the higher surface brightness regions at those contact discontinuities exhibit cooler temperatures. They infer the infall of NGC 1404 as the likely cause of this pattern. The direction of the NGC 1404 tail discussed by (Su et al., 2017a) and also seen in Fig. 5 fits this scenario, as do the hydrodynamical simulations described by Sheardown et al. (2018).
@@
The real power of eROSITA compared to all other X-ray telescopes in existence comes in through its sensitivity to large-scale low surface brightness features. Already the bottom image in Fig. 2 revealed the vast extent of warm-hot gas emission, dwarfing the emission visible in the huge XMM-Newton mosaic image shown in Fig. 5. Note that we do not expect the surface brightness fluctuations in the bottom image in Fig. 2 to be due to photon noise because of the wavelet-filtering process. Visual support for this expectation comes from comparing the eRASS:2 and eRASS3+eRASS4+eRASS5 images in Fig. 17: the two show essentially the same features despite coming from completely independent datasets. The quantitative analysis of the surface brightness profiles in different sectors follows below.
@@
To make additional use of eROSITA’s energy resolution, we show in Fig. 7 an RGB image optimized to show low surface brightness fluctuations. Hot group and cluster gas emission is revealed in this image not only through its extent but additionally through its whitish color. Figures 2 (bottom) and 7 both demonstrate that the Fornax cluster emission extends not only beyond , already unreachable by the XMM-Newton mosaic, but beyond and actually beyond even the virial radius, , in some directions. In these images, we discover finger-like emission extensions beyond in particular toward the western and southeastern directions (labeled Fingers 1, 2, and 3 in Fig. 2, bottom). These extensions appear to stretch out to well beyond , in particular Excesses 1 and 2.
@@
On the other hand, we do not see any obvious evidence in the X-ray surface brightness maps of the shock front predicted by Sheardown et al. (2018) south of NGC 1404. Below, we check quantitatively for features in the profiles in various sectors.
@@
In principle, varying diffuse emission in our Milky Way (Fig. 1) or systematic uncertainties in the relative absorption correction can result in apparent surface brightness variations. Both effects would mostly be important at the softest energies keV. Therefore, we constructed a hard band image (shown in Fig. 18) and checked if the emission features discussed above are robust. And, indeed, they are, lending confidence that they are not due to local effects.
The plots in Fig. 8 illustrate examples of surface brightness features discussed above. A bin width of was chosen for the uppermost plot as a compromise between the small sector and the relatively high surface brightness around ; for the remaining three plots in Fig. 8 we chose a bin width of as the focus lies on faint surface brightness features well beyond . The 0–360 deg profile is always shown for comparison to highlight deviations from the azimuthal average profile; its bin width always matches that of the respective comparison profile. The uppermost plot simply shows how well eROSITA can recover already known features in the inner regions; i.e., a strong drop of emission just beyond – part of the edge seen in Figs. 5 and 6. The excess emission relative to the azimuthal average is highly significant () as there are more than 10 bins above the average profile, each with a distance of several . The drop happens within about .
@@
The next plot in Fig. 8 makes clear that Excesses 1 and 2 in the west-northwest beyond discussed above have, indeed, a significantly higher surface brightness than the CXB level as there are many data points consistently above this level. Quantitatively, there are consecutive bins at least above the CXB level; i.e., the excesses are significant at the level. Furthermore, it shows that our CXB level estimate is conservative because the surface brightness of the dark blue regions is even below our CXB level estimate.
The next two plots show how the surface brightness drops abruptly and significantly at the edges of Fingers 1 and 2. Between and in the Finger 1 sector, almost all bins are well above the CXB level. Moreover, many of the bins also lie above the average profile. In particular, two consecutive bins between and lie several above the average level; they mark the edge of Finger 1, potentially consisting of a clump of gas of width up to ( kpc). Regarding Finger 2, between and there are consecutive bins above even the average profile, almost all of them by a few . Therefore, the Finger 2 excess above average is detected at significance.
@@
We could then also search for the detached bow shock predicted by Sheardown et al. (2018) to lie 450–750 kpc south of NGC 1399; i.e., roughly in the range between (393 kpc, 1.06 deg) and (816 kpc, 2.22 deg). Since it is not clear where exactly such a shock would be located and given the complexity of the two-dimensional emission in the outskirts, we started with a large sector and then looked at smaller ones in different directions. Figure 3 (bottom) shows the southern 180 deg sector. There are deviations from the average profile in the very inner region very roughly around (here and also in many of the profiles shown in Figs. 8 and 9), which are related to the spiral sloshing pattern discussed earlier. However, no obvious surface brightness jumps that might be associated with a shock are obvious in the relevant radial range. Figures 19 and 20 show several more sectors. No obvious jump is present in the range 225–315 deg. In the 240–300 deg sector, there appear to be two peaks of emission surrounding . However, consulting the surface brightness profiles of other sectors, in particular 225–270 deg and 240–270 deg, as well as Finger 3 toward the southeast in Fig. 2, it becomes clear that these are two extended (note that each individual bin covers 60′′; i.e., twice the eROSITA point spread function) clumps of gas, not related to a shock front. On the other hand, the sector 180–270 deg does show a very weak indication of a surface brightness jump between and ; however, we consider it too weak, and the overall two-dimensional emission pattern too complex, to indicate a shock.
There are several possible reasons why no obvious bow shock is detected in the predicted range. On the simulation side, the set-up might be too idealized and, for example, without sufficient substructure, or it may not cover enough plausible merger scenarios. On the observational side, the signal might be insufficient for a significant detection, or the shape of the shock may not be well captured by our sector analyses.